# ART-DECO: Arbitrary Text Guidance for 3D Detailizer Construction

### v1

@@ 1. Introduction
3D generative models are becoming increasingly powerful, enabling the creation of 3D content with ease from texts (Poole et al., 2023; Shi et al., 2024; Li et al., 2024b) or images (Liu et al., 2023; Hong et al., 2024; Xiang et al., 2024). However, artistic creation involves realizing the design vision of the artist, which often requires precise control over both the coarse structure and the local details of the generated object. Such precision cannot be fully achieved through text or image inputs alone. In addition, the artist’s creative exploration of the design space is greatly facilitated by the ability to quickly generate detailed 3D assets via structural variations.
@@
Prior works propose generative models which rely on user-defined coarse shapes to control the structure of the generated shapes, e.g., by formulating the problem as that of 3D voxel up-resolution (Chen et al., 2021, 2024a, 2023b; Shen et al., 2021a; Ren et al., 2024), while some others (Zhang et al., 2024; Hui et al., 2024) develop 3D generative models with coarse shapes as conditioning. However, as their models were trained on limited 3D shapes, they face generalizability issues and cannot produce local details of arbitrary desired styles. Some approaches (Chen et al., 2023a; Metzer et al., 2023) optimize the initial coarse shape via Score Distillation Sampling (SDS) using text-to-image diffusion models, which possess much stronger generalizability as they were trained on large image collections. Yet the structures of their generated shapes often deviate from those of the input shapes. In addition, each shape takes a significant amount of time to optimize, ranging from minutes to hours. More recently, some approaches (Chen et al., 2024b; Dong et al., 2024) have improved the structure adherence by adopting finetuned multi-view diffusion models conditioned on coarse shape inputs.
@@ 2. Related work
Figure 4. Overview of the training of our detailizer. Given a coarse voxel grid and a text prompt that describes a style, two 3D convolutional networks upsample the coarse voxels into high-resolution density and albedo fields, respectively. Multi-view images are then rendered from the density and albedo fields, and a pretrained multi-view diffusion model conditioned on the text prompt is used as a prior for Score Distillation Sampling ($\mathcal{L}_{SDS}$). The regularization loss ($\mathcal{L}_{reg}$) measures the similarity between the masks rendered from the generated shape and those from the input coarse voxel grid, thus enforcing the structure of the generated shape to be consistent with the input coarse voxels.
@@
3D generation methods went through drastic development in the recent decade. Early works mainly focus on building generative models under various 3D representations, such as voxels (Choy et al., 2016; Wu et al., 2016), point clouds (Achlioptas et al., 2018; Fan et al., 2017; Nichol et al., 2022), implicit fields (Mescheder et al., 2019; Chen and Zhang, 2019; Park et al., 2019), and polygon meshes (Nash et al., 2020; Shen et al., 2024; Siddiqui et al., 2024). These methods are trained with various generative frameworks such as variational autoencoders (Kingma and Welling, 2014), generative adversarial networks (Goodfellow et al., 2020), and denoising diffusion probabilistic models (Song et al., 2021; Ho et al., 2020). Due to data scarcity, these methods have limited generalization ability beyond the training distribution.
@@
With the success of image diffusion models (Saharia et al., 2022; Rombach et al., 2022), DreamFusion (Poole et al., 2023) and subsequent works (Wang et al., 2023; Lin et al., 2023; Melas-Kyriazi et al., 2023; Chen et al., 2023a) distill the 2D image prior in the diffusion models into 3D shapes by optimizing individual shapes with Score Distillation Sampling (SDS), therefore achieving zero-shot text-to-3D generation. Several works (Xie et al., 2024; Lorraine et al., 2023; Qian et al., 2024) also attempt to distill feed-forward text-to-3D generative models with SDS. While their generalization ability has been greatly improved, they still lack 3D understanding, which often leads to suboptimal results such as the multi-face Janus problem (Chen et al., 2023a). Zero123 (Liu et al., 2023) finetunes image diffusion models on large 3D object datasets such as Objaverse (Deitke et al., 2023), to make the diffusion model capable of generating novel-view images conditioned on a single input image and the target viewpoints. Later works including MVDream (Shi et al., 2024) and others (Shi et al., 2023; Long et al., 2024; Liu et al., 2024; Xiang et al., 2023) focus on improving multi-view consistency by generating all the views together and introducing extra information to condition the diffusion process. These multi-view images can then be used to reconstruct a 3D shape via differentiable rendering (Wang et al., 2021; Mildenhall et al., 2021; Kerbl et al., 2023) or via a feed-forward 3D shape reconstruction network (Hong et al., 2024; Li et al., 2024b).
@@
Recently, a few methods (Xiang et al., 2024; Zhang et al., 2024; Chen et al., 2024c; Ren et al., 2024; Wu et al., 2024; Li et al., 2024a; Hui et al., 2024) have been proposed to train 3D diffusion models directly from large 3D datasets, bypassing the intermediate image diffusion model. Nevertheless, the existing 3D generative models primarily focus on text or image-conditioned generation. They often lack precise control over the overall structure of the generated 3D shapes, therefore making it difficult to be incorporated into an artist’s workflow. In contrast, our method trains a feed-forward model that detailizes a coarse shape with a style specified by a text prompt, making it possible to create 3D shapes in an interactive manner where an artist can make adjustments to the coarse shape for refinement.
@@
Many traditional (Kajiya and Kay, 1989; Neyret, 1998; Zhou et al., 2006) and learning-based (Hertz et al., 2020; Berkiten et al., 2017; Yifan et al., 2022) approaches have been proposed to synthesize geometric details on coarse shapes by transferring geometric textures, which are often represented as displacement maps or geometric texture patches, from detailed shapes. In addition, neural subdivision (Liu et al., 2020) and subsequent works (Shen et al., 2021b; Chen and Zhang, 2021; Chen et al., 2023c) can also learn local geometric details in training shapes and apply them to new shapes. In another line of work, methods that are based on voxel representation can generate geometric details by replicating local voxel patches in detailed reference shapes (Chen et al., 2021, 2024a, 2023b; Sun et al., 2022). However, these methods all require detailed 3D exemplar shapes provided as style references, which limits their capability when the exemplar shapes with desired style are scarce or unavailable. To mitigate the issue of data scarcity, some recent methods (Chen et al., 2023a; Metzer et al., 2023; Gao et al., 2023; Michel et al., 2022) propose to utilize image-space supervision provided by CLIP (Radford et al., 2021) or SDS, where the style is provided by an input text. However, these methods take a long time to converge, hindering the possibility of an interactive modeling experience. Most recent works (Chen et al., 2024b; Dong et al., 2024) finetune multi-view image diffusion models conditioned on input coarse shapes, which provide faster convergence and better structure adherence. Despite these advancements, a common issue with these methods is that they can only generate shapes within the training distribution and fail when the structure of the conditioning coarse shape is uncommon. Our work addresses this issue by distilling a generative model which is designed to be generalizable and trained with increasing structure complexity.
@@ 3.1. Network architecture
We utilize 3D convolutional networks as backbone for our detailizer, as inspired by prior voxel detailization works (Chen et al., 2021, 2024a, 2023b) which showed great generalizability on arbitrary input coarse shapes. The detailed network architecture can be found in the supplementary. The input to our detailizer is a coarse occupancy voxel grid $v\in\{0,1\}^{k\times k\times k}$. The detailizer has two separate upsampling networks: $\mathbf{G}_{d}$ for generating the density field $v^{\prime}_{d}=\mathbf{G}_{d}(v)\in[0,+\infty)^{K\times K\times K}$, and $\mathbf{G}_{a}$ for the albedo field $v_{a}=\mathbf{G}_{a}(v)\in[0,1]^{K\times K\times K\times 3}$, where each voxel stores an RGB color. To enforce that the generated shape follows the structure of the input voxels, we dilate the input voxel grid $v$ by one voxel and then upsample it to $K^{3}$ resolution via nearest neighbor to obtain a binary mask $v_{m}$. We then apply the mask to the output density field $v^{\prime}_{d}$ via element-wise multiplication to obtain the final density field: $v_{d}=v^{\prime}_{d}\cdot v_{m}$. This masking step prevents the network from generating artifacts or unnecessary details far away from the structure of the input coarse shape, and allows the network to focus its capacity on generating plausible details only within the valid region $v_{m}$.
@@ 3.2. Training
To train the detailizer, we run the networks on a training set of coarse voxel grids. For each voxel grid, we generate the density field $v_{d}$ and the albedo field $v_{a}$, render multi-view images from them via volumetric rendering (Mildenhall et al., 2021), and leverage a pretrained multi-view diffusion model to provide training supervision via SDS (Poole et al., 2023; Shi et al., 2024). Formally, given the density field $v_{d}$, the albedo field $v_{a}$, the camera parameters $\mathbf{c}_{i}$ (for the i-th view), the volumetric rendering function $\mathbf{R}(\cdot)$, a text prompt $y$, and the pretrained multi-view diffusion model $\epsilon_{\theta}$, we first render multi-view images $\mathbf{x}_{i}=\mathbf{R}(v_{d},v_{a};\mathbf{c}_{i})$. Note that $\epsilon_{\theta}$ operates on multi-view images, so in each iteration, we render a set of images $\{\mathbf{x}_{i}\}_{i=1,...,N}$ from the same shape, and then we can use $\epsilon_{\theta}$ to obtain the denoised images of $\mathbf{x}_{i}$: $\{\hat{\mathbf{x}}_{i}\}=\epsilon_{\theta}(\{\mathbf{x}_{i}\}_{t};y,\{\mathbf{c}_{i}\},t)$, where $t$ is the time step and $\{\mathbf{x}_{i}\}_{t}$ are input images after adding noise $\epsilon$ for time step $t$. Now we can define our multi-view SDS loss as
@@ 4.1. Text-guided detailization
We showcase qualitative results of our method in Figure  12, and compare with other methods in Figure  14. In Figure  14, Coin3D struggles to generate accurate geometry and texture when conditioned on complex structures. This is primarily because the multi-view diffusion model that Coin3D is based on does not generalize well to uncommon structures. We provide additional results in the supplementary material to show the intermediate outputs from Coin3D and to demonstrate that Coin3D works on simple structures but fails on complex structures. ShaDDR was designed to be trained with clean, high-quality 3D shapes. Therefore, when trained with noisy shapes obtained from SDS optimization, the quality of ShaDDR’s output becomes unsatisfactory. CLAY has cleaner geometry and better structure preservation as it was trained on real 3D data. Nonetheless, it often fails to generate detailed local geometries and textures. CLAY can also deviate significantly from the structure of the input coarse shape if the structure is uncommon, see the table and animal examples in Figure  14. In comparison, our method can effectively handle both out-of-distribution coarse structures and creative text prompts. It can provide users with flexible control over the structure and the style while generating high-quality shapes. We also report quantitative comparisons in Table 1, which further demonstrates the superior performance of our method.
@@ 4.2. Ablation Study
Refer to caption(a) Baseline(b) Ours(c) Ours“a short-hairedtabby cat”1. Make thebody larger2. Reposition thetail downward3. Increase theleg spacing
Figure 7. Ablation study on the masked MVDream baseline. The baseline (a) fails to preserve the input structure (see (b) top). Our result (b) fully respects the structure, and changes its structure according to new edits (c).
@@
To show the necessity of distilling a generative neural network rather than distilling a single shape, we adopt MVDream to directly optimize a neural field while enforcing structural control by masking the neural field with the input coarse voxels. As shown in Figure 7 (a), the baseline fails to preserve the input coarse structure, as MVDream’s multi-view diffusion model lacks prior knowledge of atypical structures, such as a cat with six legs. In contrast, our model, although trained on coarse voxels of four-legged animals using the same multi-view diffusion model, demonstrates strong generalization ability to unseen structures, thanks to the inductive bias of convolutional networks. Moreover, our model can refine new edits in a single forward pass in less than a second without retraining, whereas optimization-based methods require retraining for each new structural change; see Figure 7 (c).
@@ 5. Conclusions
Refer to caption(a) Car(b) Plant“a modern SUV with a sturdy frameand a tinted windows”“an asplenium nidus plantedin a clay pot”
Figure 10. Limitations. Our current representation is not suitable for generating shapes with glossy/reflective material or shapes with thin threads/surfaces.
@@
The dense voxel representation that our method adopts enables us to take advantage of convolutional networks. However, it also limits the resolution of our generated shapes. Alternative representations and network design can be explored to address this issue. Our representation stores a fixed color in each voxel instead of view-dependent colors, therefore it cannot represent shapes with glossy or reflective surfaces, see Figure  10. Our detailizer requires re-training for each different text prompt, which can be inconvenient for fast prototyping. With sufficient compute, we believe it is possible to distill a more generic detailizer that can take both the text prompt and the coarse shape as input during inference, and generate detailed shapes at an interactive speed.

### v3

@@ 1. Introduction
3D generative models are becoming increasingly powerful, enabling the creation of 3D content with ease from texts (Poole et al., 2023; Shi et al., 2024; Li et al., 2024) or images (Liu et al., 2023; Hong et al., 2024; Xiang et al., 2025). However, artistic creation involves realizing the design vision of the artist, which often requires precise control over both the coarse structure and the local details of the generated object. Such precision cannot be fully achieved through text or image inputs alone. In addition, the artist’s creative exploration of the design space is greatly facilitated by the ability to quickly generate detailed 3D assets via structural variations.
@@
Prior works propose generative models which rely on user-defined coarse shapes to control the structure of the generated shapes, e.g., by formulating the problem as that of 3D voxel up-resolution (Chen et al., 2021, 2024a, 2023b; Shen et al., 2021; Ren et al., 2024), while some others (Zhang et al., 2024; Hui et al., 2024) develop 3D generative models with coarse shapes as conditioning. However, as their models were trained on limited 3D shapes, they face generalizability issues and cannot produce local details of arbitrary desired styles. Some approaches (Chen et al., 2023a; Metzer et al., 2023) optimize the initial coarse shape via Score Distillation Sampling (SDS) using text-to-image diffusion models, which possess much stronger generalizability as they were trained on large image collections. Yet the structures of their generated shapes often deviate from those of the input shapes. In addition, each shape takes a significant amount of time to optimize, ranging from minutes to hours. More recently, some approaches (Chen et al., 2024b; Dong et al., 2024) have improved the structure adherence by adopting finetuned multi-view diffusion models conditioned on coarse shape inputs.
@@ 2. Related work
Figure 4. Overview of the training of our detailizer. Given a coarse voxel grid and a text prompt that describes a style, two 3D convolutional networks upsample the coarse voxels into high-resolution density and albedo fields, respectively. Multi-view images are then rendered from the density and albedo fields, and a pretrained multi-view diffusion model conditioned on the text prompt is used as a prior for Score Distillation Sampling ($\mathcal{L}_{SDS}$). The regularization loss ($\mathcal{L}_{reg}$) enforces structural consistency by measuring the similarity between the masks rendered from the generated shape and those from the input coarse voxel grid. During inference time, the detailizer upsamples a coarse voxel grid into high-resolution fields in a single feedforward pass, without SDS optimization.
@@
3D generation methods went through drastic development in the recent decade. Early works mainly focus on building generative models under various 3D representations, such as voxels (Choy et al., 2016; Wu et al., 2016), point clouds (Achlioptas et al., 2018; Fan et al., 2017; Nichol et al., 2022), implicit fields (Mescheder et al., 2019; Chen and Zhang, 2019; Park et al., 2019), and polygon meshes (Nash et al., 2020; Shen et al., 2024; Siddiqui et al., 2024). These methods are trained with various generative frameworks such as variational autoencoders (Kingma and Welling, 2014), generative adversarial networks (Goodfellow et al., 2014), and denoising diffusion probabilistic models (Song et al., 2021; Ho et al., 2020). Due to data scarcity, these methods have limited generalization ability beyond the training distribution.
@@
With the success of image diffusion models (Saharia et al., 2022; Rombach et al., 2022), DreamFusion (Poole et al., 2023) and subsequent works (Wang et al., 2023; Lin et al., 2023; Melas-Kyriazi et al., 2023; Chen et al., 2023a) distill the 2D image prior in the diffusion models into 3D shapes by optimizing individual shapes with Score Distillation Sampling (SDS), therefore achieving zero-shot text-to-3D generation. Several works (Xie et al., 2024; Lorraine et al., 2023; Qian et al., 2024) also attempt to distill feed-forward text-to-3D generative models with SDS. While their generalization ability has been greatly improved, they still lack 3D understanding, which often leads to suboptimal results such as the multi-face Janus problem (Chen et al., 2023a). Zero123 (Liu et al., 2023) finetunes image diffusion models on large 3D object datasets such as Objaverse (Deitke et al., 2023), to make the diffusion model capable of generating novel-view images conditioned on a single input image and the target viewpoints. Later works including MVDream (Shi et al., 2024) and others (Shi et al., 2023; Long et al., 2024; Liu et al., 2024a; Xiang et al., 2023) focus on improving multi-view consistency by generating all the views together and introducing extra information to condition the diffusion process. These multi-view images can then be used to reconstruct a 3D shape via differentiable rendering (Wang et al., 2021; Mildenhall et al., 2020; Kerbl et al., 2023) or via a feed-forward 3D shape reconstruction network (Hong et al., 2024; Li et al., 2024).
@@
Recently, a few methods (Xiang et al., 2025; Zhang et al., 2024; Chen et al., 2025; Ren et al., 2024; Wu et al., 2024; Li et al., 2025; Hui et al., 2024) have been proposed to train 3D diffusion models directly from large 3D datasets, bypassing the intermediate image diffusion model. Nevertheless, the existing 3D generative models primarily focus on text or image-conditioned generation. They often lack precise control over the overall structure of the generated 3D shapes, therefore making it difficult to be incorporated into an artist’s workflow. In contrast, our method trains a feed-forward model that detailizes a coarse shape with a style specified by a text prompt, making it possible to create 3D shapes in an interactive manner where an artist can make adjustments to the coarse shape for refinement.
@@
Many traditional (Kajiya and Kay, 1989; Neyret, 1998; Zhou et al., 2006) and learning-based (Hertz et al., 2020; Berkiten et al., 2017; Yifan et al., 2022) approaches have been proposed to synthesize geometric details on coarse shapes by transferring geometric textures, which are often represented as displacement maps or geometric texture patches, from detailed shapes. In addition, neural subdivision (Liu et al., 2020) and subsequent works (Shen et al., 2021; Chen and Zhang, 2021; Chen et al., 2023c) can learn local geometric details in training shapes and apply them to new shapes. In another line of work, methods that are based on voxel representation can generate geometric details by replicating local voxel patches in detailed reference shapes (Chen et al., 2021, 2024a, 2023b; Sun et al., 2022). However, these methods all require detailed 3D exemplar shapes provided as style references, which limits their capability when the exemplar shapes with desired style are scarce or unavailable. To mitigate the issue of data scarcity, some recent methods (Chen et al., 2023a; Metzer et al., 2023; Gao et al., 2023; Michel et al., 2022) propose to utilize image-space supervision provided by CLIP (Radford et al., 2021) or SDS, where the style is provided by an input text. However, these methods take a long time to converge, hindering the possibility of an interactive modeling experience. Most recent works (Chen et al., 2024b; Dong et al., 2024; Liu et al., 2024b) leverage multi-view image diffusion models conditioned on input coarse shapes, which provide faster convergence and better structure adherence. Despite these advancements, a common issue with these methods is that they can only generate shapes within the training distribution and fail when the structure of the conditioning coarse shape is uncommon. Our work addresses this issue by distilling a generative model which is designed to be generalizable and trained with increasing structure complexity.
@@ 3.1. Network architecture
We utilize 3D convolutional networks as backbone for our detailizer, as inspired by prior voxel detailization works (Chen et al., 2021, 2024a, 2023b) which showed great generalizability on arbitrary input coarse shapes. The detailed network architecture can be found in the supplementary. The input to our detailizer is a coarse occupancy voxel grid $v\in\{0,1\}^{k\times k\times k}$. The detailizer has two separate upsampling networks: $\mathbf{G}_{d}$ for generating the density field $v^{\prime}_{d}=\mathbf{G}_{d}(v)\in[0,+\infty)^{K\times K\times K}$, and $\mathbf{G}_{a}$ for the albedo field $v_{a}=\mathbf{G}_{a}(v)\in[0,1]^{K\times K\times K\times 3}$, where each voxel stores an RGB color. To enforce that the generated shape follows the structure of the input voxels, we dilate the input voxel grid $v$ by one voxel (Chen et al., 2023b) and then upsample it to $K^{3}$ resolution via nearest neighbor to obtain a binary mask $v_{m}$. We then apply the mask to the output density field $v^{\prime}_{d}$ via element-wise multiplication to obtain the final density field: $v_{d}=v^{\prime}_{d}\cdot v_{m}$. This masking step prevents the network from generating artifacts or unnecessary details far away from the structure of the input coarse shape, and allows the network to focus its capacity on generating plausible details only within the valid region $v_{m}$.
@@ 3.2. Training
To train the detailizer, we run the networks on a training set of coarse voxel grids. For each voxel grid, we generate the density field $v_{d}$ and the albedo field $v_{a}$, render multi-view images from them via volumetric rendering (Mildenhall et al., 2020), and leverage a pretrained multi-view diffusion model to provide training supervision via SDS (Poole et al., 2023; Shi et al., 2024). Formally, given the density field $v_{d}$, the albedo field $v_{a}$, the camera parameters $\mathbf{c}_{i}$ (for the i-th view), the volumetric rendering function $\mathbf{R}(\cdot)$, a text prompt $y$, and the pretrained multi-view diffusion model $\epsilon_{\theta}$, we first render multi-view images $\mathbf{x}_{i}=\mathbf{R}(v_{d},v_{a};\mathbf{c}_{i})$. Note that $\epsilon_{\theta}$ operates on multi-view images, so in each iteration, we render a set of images $\{\mathbf{x}_{i}\}_{i=1,...,N}$ from the same shape, and then we can use $\epsilon_{\theta}$ to obtain the denoised images of $\mathbf{x}_{i}$: $\{\hat{\mathbf{x}}_{i}\}=\epsilon_{\theta}(\{\mathbf{x}_{i}\}_{t};y,\{\mathbf{c}_{i}\},t)$, where $t$ is the time step and $\{\mathbf{x}_{i}\}_{t}$ are input images after adding noise $\epsilon$ for time step $t$. Now we can define our multi-view SDS loss as
@@ 4.1. Text-guided detailization
We showcase qualitative results of our method in Figure  12, and compare with other methods in Figure  14. In Figure  14, Coin3D struggles to generate accurate geometry and texture when conditioned on complex structures. This is primarily because the multi-view diffusion model that Coin3D is based on does not generalize well to uncommon structures. We provide additional results in the supplementary material to show the intermediate outputs from Coin3D and to demonstrate that Coin3D works on simple structures but fails on complex structures. ShaDDR was designed to be trained with clean, high-quality 3D shapes. Therefore, when trained with noisy shapes obtained from SDS optimization, the quality of ShaDDR’s output becomes unsatisfactory. CLAY has cleaner geometry and better structure preservation as it was trained on real 3D data. Nonetheless, it often fails to generate detailed local geometries and textures. CLAY can also deviate significantly from the structure of the input coarse shape if the structure is uncommon, see the table and animal examples in Figure  14. In comparison, our method can effectively handle both out-of-distribution coarse structures and creative text prompts. It can provide users with flexible control over the structure and the style while generating high-quality shapes. We report quantitative comparisons in Table 1, which further demonstrates the superior performance of our method. We also provide a user study in the supplementary to show that participants clearly prefer our results over alternatives.
@@ 4.2. Ablation Study
Refer to caption(a) Input coarse voxels(b) Baseline(c) Ours“a short-hairedtabby cat”
Figure 7. Ablation study on the masked MVDream baseline. The baseline (b) fails to preserve the input structure, while our method (c) fully respects it.
@@
To show the necessity of distilling a generative neural network rather than distilling a single shape, we adopt MVDream to directly optimize a neural field while enforcing structural control by masking the neural field with the input coarse voxels. As shown in Figure 7 (b), the baseline fails to preserve the input coarse structure, as MVDream’s multi-view diffusion model lacks prior knowledge of atypical structures, such as a cat with six legs. In contrast, our model, although trained on coarse voxels of four-legged animals using the same multi-view diffusion model, demonstrates strong generalization ability to unseen structures, thanks to the inductive bias of convolutional networks.
@@ 5. Conclusions
Refer to caption(a) Table(b) Car(c) Plant“a stone table with aheavy base and texture”“a modern SUV with a sturdyframe and tinted windows”“an asplenium nidusplanted in a clay pot”
Figure 10. Limitations. Our current representation is constrained in voxel resolution and is not suitable for generating shapes with glossy/reflective material or thin surfaces/threads. (a) left shows the geometry of the shape.
@@
Limitation.
The dense voxel representation that our method adopts enables us to take advantage of convolutional networks. However, it limits the resolution of our generated shapes; the surfaces can appear blocky or jagged due to voxel artifacts, causing local details and smooth curves to be less faithfully presented. It also struggles with thin structures such as plant leaves; see Figure 10 (a) and (c). Alternative representations and network design can be explored to address this issue. Our representation stores a fixed color in each voxel instead of view-dependent colors, therefore it cannot represent shapes with glossy or reflective surfaces, see Figure 10 (b). Our detailizer requires re-training for each different text prompt, which can be inconvenient for fast prototyping. Lastly, our method requires a set of category-specific coarse shapes for training each new style, which limits its applicability in domains where such data does not exist, and makes it less feasible to scale across many styles without additional human effort. With sufficient compute, we believe it is possible to distill a more generic detailizer that can take both the text prompt and the coarse shape as input during inference, and generate detailed shapes at an interactive speed.